Find the area of the following circles, given that:
Radius = 14 mm (Take π = 22/7)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given, radius of circle = 14 mm
Then,
Area of the circle =πr^2
= 22/7 × 14^2
= 22/7 × 196
= 22 × 28
= 616 mm^2

Thanks